OpenStudy (anonymous):

To write 93,000,000.0 in scientific notation, what is the mantissa and the exponent? (Enter the mantissa to one decimal place.)

OpenStudy (shamim):

In scientific notation 100=1*10^2 Right?

OpenStudy (shamim):

1000=1*10^3

OpenStudy (shamim):

10000=1*10^4

OpenStudy (michele_laino):

hint: we can rewrite your number as below: \[\Large 9.3 \times {10^7}\]